### v2

# CRS-Diff: Controllable Generative Remote Sensing Foundation Model

Abstract
The emergence of diffusion models has revolutionized the field of image generation, providing new methods for creating high-quality, high-resolution images across various applications. However, the potential of these models for generating domain-specific images, particularly remote sensing (RS) images, remains largely untapped. RS images that are notable for their high resolution, extensive coverage, and rich information content, bring new challenges that general diffusion models may not adequately address. This paper proposes CRS-Diff, a pioneering diffusion modeling framework specifically tailored for generating remote sensing imagery, leveraging the inherent advantages of diffusion models while integrating advanced control mechanisms to ensure that the imagery is not only visually clear but also enriched with geographic and temporal information. The model integrates global and local control inputs, enabling precise combinations of generation conditions to refine the generation process. A comprehensive evaluation of CRS-Diff has demonstrated its superior capability to generate RS imagery both in a single condition and multiple conditions compared with previous methods in terms of image quality and diversity. The code is available at https://github.com/Sonettoo/CRS-Diff.
Keywords:
Remote sensing image Diffusion model Controllable generation
Figure 1: Visualisation results of our proposed CRS-Diff. (a) Single Image generation: The first line shows the examples of generated RS images based on only text. (b) Single-conditional image generation: The second line shows the generation results under the control of one additional condition. (c) Multi-conditional image generation: The last line shows the generation results under the control of multiple conditions.
1 Introduction
Diffusion models [10, 36] are a class of probabilistic generative models that turn noise into a representative data sample. Recently, image generation based on diffusion -model [20, 5, 24, 29, 2, 6, 35, 22] has emerged as a hot research topic since the generated images exhibit high quality, e.g., generating realistic images [20, 5, 24, 29], transforming art styles [39, 48], image super-resolution [41, 14, 31], video generation [9], etc. However, most existing diffusion models focus primarily on general image generation, with insufficient exploration in generating specific types of images, such as remote sensing (RS) images.
RS images differ from traditional RGB images in several aspects, particularly in terms of resolution, coverage area, and informational richness. Covering vast geographical areas, these images typically contain a wealth of surface and atmospheric data. These unique characteristics of RS images thus bring a new challenge for RS image generation, i.e., relying solely on textual control is often inadequate for generating RS images of high quality that meet specific requirements. Thus, more controlled conditions for RS image generation are needed to be explored. Currently, there have been several controlled RS image generation models [4, 13]. For example, [4] shows that the pre-trained diffusion models can be conditioned on cartographic data to generate realistic satellite images. DiffusionSat [13] incorporates the associated metadata such as geolocation as conditioning information. However, these models overlook the use of image-related features as control signals. Additionally, these models lack the flexibility of combining any conditions to generate images.
To address these issues, in this paper, we propose CRS-Diff, i.e., a controllable generative remote sensing foundation model. Specifically, a base diffusion model is first trained tailored for the RS domain based on the Stable Diffusion (SD) model[29], which is capable of converting high-precision textual descriptions into RS images as shown in fig. 1. Building upon this, we integrated two additional control signals into the diffusion model to enable the controlled generation of RS images. These two control signals manage the adjustment of the global and local information of the images, which includes seven additional image control conditions (e.g. semantic segmentation masks, road maps, sketch, etc) as shown in Fig.fig. 1(b). Furthermore, incorporating metadata into the timestep embeddings introduces spatial-temporal information, ensuring that the generated RS images are not only visually realistic but also accurately reflect specific geographical and temporal information. Finally, for each control condition, we can manually set whether to choose that condition, thus making our method flexible to combine any conditions to generate images. In summary, the contributions are threefold:
  1. 1.
    We propose a new controllable generative RS foundation model (CRS-Diff).
  2. 2.
    We introduce a new conditional control mechanism to broaden the image generation space and achieve the effective integration of multiple control information.
  3. 3.
    Experimental results have demonstrated the superiority of our CRS-Diff in generating RS imagery that adheres to specific conditions and surpasses previous RS image generation methods in terms of image quality and diversity.
2 Related work
2.1 Text-to-Image Generation
Text-to-image generation has attracted significant attention in the multimodal field, to generate high-definition images corresponding to given textual descriptions. Early research was focused primarily on GANs[27, 28, 46], with text-conditional GANs emerging as the pioneering end-to-end differential architectures from character to pixel level. For instance, Reed et al. introduced the generative adversarial what-where network, capable of generating 128×128 pixel images. Additionally, Reed et al. developed StackGAN, employing stacked generators to produce clear 256 × 256 pixel remote sensing images. However, these models encounter two main challenges: training instability and limited generalization to open-domain scenes. In addition to GAN-based methods, recent studies have transitioned towards autoregressive models or diffusion models, such as DALL-E[24]. Yong et al. [45] utilized modern Hopfield layers[3, 33] for hierarchical prototype learning[16] in text and image embeddings, to extract the most representative prototypes and implement a coarse-to-fine learning strategy, thereby enhancing the realism of generated remote sensing images.
Diffusion Models[10] are generative models that generate new images by gradually transforming an image from a Gaussian noise state to a target image. This process is composed of two main steps: the forward diffusion process and the reverse generation process. Compared to autoregressive models, diffusion models excel in generating more realistic images through a gradual denoising process. Numerous studies have since focused on enhancing the diffusion model, DALLE-2[24] enhances textual guidance capabilities through integration with the CLIP model, while GLIDE[20] explores diverse guidance methodologies. Conversely, Stable Diffusion(SD)[29] augments training data leveraging the capabilities of the diffusion model, thus improving the generation results.
2.2 Controlled diffusion models
The application of Controlled Diffusion Models (CDM)[47, 12, 19, 13, 50, 15] in the field of Text-to-Image (T2I) generation aims to enable users to precisely dictate the content of generated images. Although traditional text-to-image (T2I) diffusion models are capable of generating images from text descriptions, users often experience limited control over the final output. Controlled diffusion models enable users to specify additional generative details, including style, color, and object positioning, through the introduction of enhanced controllable parameters or mechanisms. For example, ControlNet[47], GLIGEN[15] and T2I-Adapter[19] all incorporate additional control networks or control signals on top of the SD weights to enable integrated control of multiple conditions to reduce training costs. Composer trains a large diffusion model from scratch through a new generative paradigm that allows for flexibility in constructing generative conditions, improving controllability, and achieve just the right results.
In particular, ControlNet[47] can be used in conjunction with diffusion models such as DDPM and LDM to augment the controllability and diversity of the generated images. By introducing additional control signals or conditions, such as textual descriptions, image attribute labels, etc., Based on ControlNet, Zhao et al. [50] proposed that Uni-ControlNet is capable of accepting a wide array of additional control signals or combinations of conditions. By merely fine-tuning the adapters while keeping the original SD model unchanged, Uni-ControlNet significantly reduces the training cost, necessitating only two additional adapters for effective outcomes.
In the field of remote sensing, recent work by Espinosa et al. [4] proposed that pre-trained diffusion models can be conditioned on cartographic data to generate realistic satellite images. DiffusionSat[13], has demonstrated the capability to generate high-resolution satellite data utilizing numerical metadata and textual captions. This approach can be extended to address inverse problems, such as multi-spectral input super-resolution. However, DiffusionSat lacks the capability to incorporate input conditions of additional dimensions and enhanced textual control. In contrast, the introduction of an additional control network enables our model to achieve composite control generation under arbitrary conditions and synthesize images with higher precision, attributable to our sophisticated training strategy.
3 Method
We introduce the diffusion model into the field of remote sensing image generation, aiming to enhance generic image generation capabilities for producing more realistic remote sensing images, subsequently introducing an optimized multi-conditional control mechanism that leverages text, image, and other multidimensional information to guide precise image generation and yield high-precision remote sensing images. The construction of the model consists of two steps: initially, text-image pairs are utilized to train the generative diffusion model weights for remote sensing images, building upon the traditional SD framework, and then the combination of multiple conditions (local conditions and global conditions) is implemented through a conditional control network. We will detail the two-stage training process for CRS-Diff, alongside illustrating the implementation of separate decompositions and combinations of training data.
3.1 Single Image Generation
Initially, we employed a diffusion model-based approach for text-to-image generation. This process involves utilizing a frozen variational autoencoder (VAE) encoder and decoder. The purpose is to convert each image $x\in\mathbb{R}^{C\times H\times W}$ into its corresponding latent space variable $z$, thus circumventing the direct learning of the original image’s conditional data distribution given the text conditions $p(x|\tau)$, and instead focuses on learning the feature distribution of the mapped image feature vector $p(z|\tau)$. Text description corresponding to an image is encoded by a CLIP model[23], which subsequently guides the image generation during the denoising process via a cross-attention mechanism[38]. Additionally, a Classifier-free Guidance (CFG) mechanism[11] is introduced:
$\hat{\theta}(z_{t},c)=\omega\theta(z_{t},c)+(1-\omega)\theta(z_{t}),$ (1)
where $z_{t}=a_{t}x_{0}+\sigma_{t}$ and $\omega$ denotes the bootstrap weight. $\hat{\theta}(z_{t},c)$ is the output of the CFG mechanism. It is a weighted sum of the class-conditional output $\theta(z_{t},c)$ and the unconditional output $\theta(z_{t})$.
Textual information serves as the sole guiding factor in this process. Simultaneously, we employ the pretrained CLIP ViT-L/14@336px model[29], fine-tuned on the RSICD remote sensing image dataset, to amplify the effect of textual guidance. Under the original CFG framework, conditional noise prediction $\theta(x_{t},c)$ is solely dependent on the model’s processing of a given condition (c) (e.g., text description). With the introduction of the CLIP bootstrap, this conditional prediction becomes further influenced by the similarity loss between the image and text as computed by the CLIP model. This implies that the calculation of $\theta(x_{t},c)$ incorporates considerations for better aligning the resulting image with the textual description.
Our text-to-image generation methodology initially encodes an image $x\in\mathbb{R}^{C\times H\times W}$ utilizing a static VAE, which is then converted into a latent representation $z=E(x)\in\mathbb{R}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$. Subsequently, Gaussian noise is added to the latent image features to produce a noisy latent representation $z_{t}=\alpha_{t}z+\sigma_{t}\varepsilon$, where $\varepsilon$ represents the Gaussian noise component, and $\alpha_{t}$ and $\sigma_{t}$ are coefficients modulating the noise intensity. The text caption $\tau$ is encoded by a CLIP model $T_{\theta}(\tau)$, generating the text embedding $\tau^{\prime}$. This is accomplished through a denoising model $\epsilon_{\theta}(z_{t};\tau^{\prime},c)$. Finally, the denoised potential representation is up-sampled to the original image resolution using the VAE decoder, completing the image generation process. In this process, we use a sampling algorithm, Denoising Diffusion Implicit Models (DDIM), which is used to speed up the sampling process of diffusion models.
Additionally, the weights of the encoder $E$, decoder $D$, CLIP text encoder $T_{\theta}$, and denoising model $\epsilon_{\theta}$ were inherited from the SD 1.5 version. During the training process, only the denoising model $\epsilon_{\theta}$ is updated to accelerate training convergence and exploit the rich semantic information in the SD model.
3.2 Image Decoupling
To address the issue of insufficient training data, the original image data is decoupled into the corresponding feature condition data using supplementary network structures, thus constructing a large-scale combined condition dataset that also functions as an interface during the inference process, and subsequently introducing nine conditions for formal model training.
Caption: Under conditions where caption data is available, such as the RSICD dataset[18], we directly utilize the corresponding captions of remote sensing images. For other datasets, such as fMoW dataset[1], we leverage the category information of remote sensing images, to synthesize captions.
HED: A pre-trained deep neural network[42] is employed to predict edges and object boundaries directly from the original image, thereby capturing high-level object boundary information and low-level details.
MLSD: We use a pre-trained transformer-based model[44] to detect straight line segments in remote sensing images.
Depthmap: We use a pre-trained depth estimation model[26] to extract the depth map of the image, which approximates the layout of the image and aids in enhancing the model’s understanding of the remote sensing image’s semantics.
Sketch: Edge detection models[34] are applied to extract sketches from an image, focusing on the local details while conveying limited semantics.
Road Map: In certain remote sensing images, greater emphasis is placed on road information, leading to the introduction of a pre-trained Separable Graph Convolutional Network (SGCN)[51] aimed at road extraction to yield single-channel road data.
Segmentation Mask: we employ UNetFormer[40], a network specializing in remote sensing images, to extract semantic information and produce masks segmented into eight categories.
Content: For content, the given image is considered as control information. Utilizing the Image Encoder in the pretrained CLIP ViT-L/14@336px[23] model, the image is transformed into feature encoding, obtaining a global embedding. This approach offers more relevant guiding conditions than text descriptions alone.
Metadata: In the processing of remote sensing images, the inclusion of additional metadata, including temporal (year, month, day) and spatial (Ground Sample Distance, latitude, longitude, cloud cover) information, is considered. This information is first quantified and categorized, serving as input for the traditional diffusion model’s category guidance[29], through the labeling of these categories. Furthermore, these categories are transformed into textual descriptions with semantic information, acting as a weak textual control condition.
3.3 Multi-conditional control
Additional conditional control modules are added based on the backbone structure outlined in the previous section to reconstruct the original image from optional decoupled image representation conditions. The known conditions are categorized into two types: global control and local control, with the corresponding conditional network constructed accordingly. These are then integrated with ControlNet’s control strategy[47] to enable composite control over arbitrary conditions.
Figure 2: The detailed depiction of CRS-Diff, where the Local Control structure (pink area) extracts multiple conditional image information via a feature extraction network and channels it into a trainable SD Encoder Block, subsequently integrating all information within another, frozen SD Encoder Block (blue area).
Global Control: We aim to establish a joint conditioning mechanism encompassing Caption, Content, and Metadatafig. 2. The Caption is maintained as the primary bootstrap condition, obtaining word embedding representations via a specialized CLIP text encoder, denoted as $y_{\text{c}}$. Additionally, the processed metadata and content are converted into embedding vectors through the CLIP text encoder and image encoder, respectively denoted as $y_{\text{m}}$ and $y_{\text{i}}$. Subsequently, these are transformed into eight additional feature tokens and concatenated with the caption encoded tokens via a Feed Forward Network (FFN), represented as:
$y_{\text{input}}^{\prime}=\text{Concat}\left(y_{\text{c}},w_{m}\text{FFN}_{m}(y_{\text{m}}),w_{i}\text{FFN}_{i}(y_{\text{i}})\right)$ (2)
where $w_{m}$ and $w_{i}$ are the weights applied to the outputs of the FFN for $y_{\text{m}}$ and $y_{\text{i}}$, respectively. This network then integrates these encodings with the caption’s word embedding, replacing the original text encoding as the Key and Value in the cross-attention layer.
Local Control: All decoupled image conditions, including sketches, segmentation masks, depth maps, road maps, and MLSD, are utilized as local control information. As shown in fig. 2, this part of CRS-Diff is based on the ControlNet, with the SD weights initially fixed to replicate the encoders and the structure and weights of the intermediate blocks. Notably, the Feature Extractor, comprising stacked convolutional neural networks, is designed to process the multiple local control information to match the dimensions of the noisy image. This processed information is then passed through a zero convolution layer, summed with the input noise, and forwarded to the replicated encoder, facilitating control over the image conditions.
Given a set of image conditions $\{c_{i}\}_{i=1}^{n}$, the local control information processing can be represented as a function $\mathcal{F}$, which maps the set of conditions to a feature space that aligns with the input noise dimensions. This mapping is formally defined as :$\mathcal{F}:\{c_{i}\}\rightarrow\mathbb{R}^{d\times d}$, where $d$ represents the dimensionality of the input noise latent. The transformation leverages a series of convolutional and pooling layers to effectively capture the spatial hierarchies and semantic features of the control information, ensuring that the generated features are representative of the underlying conditions.
We integrate information from various scales through self-attention, where the transformed control information is channelled through the FDN, merged with the input noise. Here, we harmonize denoised images and local conditions via Feature Denormalization (FDN)[21] to accommodate replicated encoders of varying resolutions[50]. The merged latent space variables are subsequently introduced into a frozen encoder block (or intermediate block).By integrating local control dynamics into the generation process, we ensure the model generates images that are not only diverse but also adhere to the specified conditions with high fidelity.
3.4 Training Strategy
Initial training is conducted using SD1.5 weights on a recognized text-to-image remote sensing dataset, aiming to develop a high-precision diffusion model for text-to-image generation that serves as the backbone of the ControlNet structure. This foundation enables the model to accurately guide the denoising process, leveraging a combination of multiple conditions through joint training. For both local and global control conditions, individual conditions are omitted with an independent probability of 0.5, and all conditions with a joint probability of 0.1, in accordance with Classifier-free Guidance. The dropout probability of certain conditions will be adjusted during experiments based on performance. Each metadata element is treated as a distinct bootstrap condition, with single or multiple conditions being omitted at certain probabilities, enabling the model to learn a broader array of condition combinations.
4 Experiments
4.1 Datasets
During the training of the model, we used multiple datasets:
RSICD dataset[18]. Designed specifically for image captioning in remote sensing imagery, the RSICD contains 10,921 aerial remote sensing images accompanied by captions in natural language, sized at $224\times 224$.
fMoW dataset.[1] A large-scale remote sensing image dataset, the Functional Map of the World (fMoW), includes spatio-temporal and category information for each image. Notably, only the RGB images, sized at $224\times 224$, are used, with relevant metadata extracted from a total of 110,000 images for training the multi-conditional control model.
Million-AID dataset[17]. A benchmark dataset for remote sensing scene classification, Million-AID contains millions of instances, featuring 51 scene categories with 2,000 to 45,000 images per category.
4.2 Implementation details
In the initial backbone model training phase, the following steps are taken: The RSICD dataset is fine-tuned over 10 epochs using a UNet network and the AdamW optimizer with a learning rate of $1\times 10^{-5}$. The input images are resized to $512\times 512$, with the model’s parameters totaling approximately 0.9 billion. For the sampling process, DDPM is utilized, setting the number of time steps to 100 and the classifier-free bootstrap scale to 7.5.
During the conditional control phase: Original images from the fMoW and Million-AID datasets are organized into 200,000 text-image pairs. These images are then segmented into multiple conditional representations through annotator networks. This process involves randomly combining single or multiple conditional information, including road maps, MLSD, content, and extracted raw captions, metadata, and fine-tuning the conditional control network over 5 epochs. The AdamW optimizer, with a learning rate of $1\times 10^{-5}$, is used throughout this process, resizing both the input image and the local conditional graph to $512\times 512$. Experiments are conducted on NVIDIA GeForce RTX 4090 and NVIDIA RTX A100 GPUs, with a batch size of 8.
4.3 Evaluation metrics
We utilize four metrics to assess the effectiveness of image generation under the single text condition, namely the Inception Score (IS)[32], the Fréchet Inception Distance (FID)[8], the CLIP Score[23], and the Overall Accuracy (OA)[45] for zero-shot classification. We evaluate the performance of our proposed method and the baselines under identical settings.
Notably, we utilize the zero-sample classification Overall Accuracy (OA) to evaluate the ability of the generative model to generalize across unseen categories. The computation of the OA for zero-sample classification is summarized in the following steps:
  1. 1.
    Train a classifier (Resnet network[7]) utilizing the images produced by the generative model as training data.
  2. 2.
    Test the accuracy of this classifier in categorizing images within a set of real, yet unseen categories encountered during the training phase. The formula is:
$OA=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}(y_{i}=\hat{y}_{i})$ (3)
where $y_{i}$ is the true label, $\hat{y}_{i}$ is the predicted label, $\mathbf{1}$ denotes the indicator function, and $n$ is the number of samples in the test set.
Figure 3: A schematic representation of remote sensing images is produced using different text-to-image generation methods based on text descriptions from the RSICD test set.
Table 1: Comparisons results between our CRS-Diff and other methods in terms of four evaluation metrics (i.e., Inception score, FID score, CLIP score, and Zero-Shot classification OA on the RSICD test set). The best results are in bold
Method IS $\uparrow$ FID $\downarrow$ CLIP Score ($\times$100) $\uparrow$ Zero-Shot classification OA $\uparrow$
Attn-GAN[43] 11.71 95.81 20.19 32.56
DAE-GAN[30] 7.71 93.15 19.69 29.74
StrucGAN[49] 5.84
DF-GAN[37] 9.51 109.41 19.76 51.99
Lafite[52] 10.70 74.11 22.52 49.37
DALL-E[25] 2.59 191.93 20.13 28.59
Txt2Img-MHN(VQVAE)[45] 3.51 175.36 21.35 41.46
Txt2Img-MHN(VQGAN) 5.99 102.44 20.27 65.72
CRS-Diff(ours) 18.39 50.72 20.33 67.38
4.4 Comparison and Analysis
4.4.1 Single Image generation.
We trained CRS-Diff on the RSICD dataset using solely text as the initial input and compared it with recent SOTA methods.
Qualitative analysis. In fig. 3, examples of generation using various methods, such as CRS-Diff, Txt2Img-MHN (including VQVAE and VQGAN)[45], are illustrated. Notably, our proposed CRS-Diff demonstrates the capability to generate clearer and more realistic images. For instance, when confronted with complex textual descriptions, exemplified by the phrase on the left side of the fifth line: "a square lawn and a half round lawn consist the square which is surrounded by the forest", the model proficiently deciphers the semantic content. Moreover, the model adeptly identifies the semantic information correlated with the textual descriptions, accurately reflecting this in the results. It also excels at synthesizing more appropriate images in response to shape descriptions like "square" and "long." Furthermore, CRS-Diff distinctly grasps the concept of quantity, a feature intuitively evident in its handling of numerical descriptors like "some", "many", or "two" (as seen on the left side of the first line). Additionally, our model can accurately simulate real lighting conditions, and coordinate elements such as color and texture.
Quantitative analysis: We compared CRS-Diff with Attn-GAN[43], DAE-GAN[30], StrucGAN[49], DF-GAN[37], Lafite[52], DALL-E[25], Txt2Img-MHN (including VQVAE and VQGAN)[45]. The generated results were quantitatively analysed using the RSICD validation set, employing four evaluation metrics: zero-shot classification OA, initial score, CLIP score, and FID score, as delineated in table 1. Our method surpassed the baseline in three metrics and achieved third place in CLIP score. Intriguingly, the performance in CLIP score was not as high as anticipated, possibly due to the specificities of the CLIP model employed in our evaluation. Nevertheless, the proposed CRS-Diff demonstrates excellent performance in controllability and generation quality, meeting the demands of practical applications like urban planning and laying a foundation for future research in controlled generation.
Figure 4: Qualitative results of controlled images generated from a single condition. All of these conditions, with the exception of the Content condition, were accompanied by textual descriptions for participation in the bootstrapping process.
4.4.2 Compositional Image Generation.
In addition to the textual description, more bootstrap conditions were added to further guide the model towards generating more refined results. Single-conditional image generation. Figure 4 illustrates the image generation results achieved under the control of a single condition. To mitigate the risk of conflicts between the Content condition and the semantics of the textual description, resulting in inaccurate results, textual bootstrapping was added to all conditions except the Content condition. Among these, HED Boundary and Sketch contribute to intuitive image control by restricting the boundary and contour information of the generated image, making them suitable for mapping simulations; while Segmentation Mask, Road Map, and Content offer richer semantic information. Metadata control appears significantly more challenging, and efforts were made to select among the more prominent attributes, such as month and Cloud Cover. Under appropriate conditions, this is reflected. This information is effectively interpreted by CRS-Diff when combined with the appropriate prompts, influencing the generated results. Semantic segmentation and road detection are associated with two sub-tasks in the remote sensing image domain, thereby broadening the usability of CRS-Diff in the remote sensing domain. In contrast, Depthmap and MLSD provide limited information; Depthmap is capable of responding to the light and dark changes in remote sensing images under specific conditions, whereas MLSD conditions are more apt for architectural design simulation tasks and offer effective guidance for image synthesis.
Figure 5: More visual results of CRS-Diff under multiple condition control. The first three lines display the results of random combinations of some conditions and text descriptions, while the last two lines illustrate the results of the experiment after removing the text descriptions.
Multi-conditional image generation. We conducted tests under multiple conditions, but it is worth noting that conflicts may arise among these conditions, and in the context of remote sensing images, even minor distortions can render the entire image meaningless, distinguishing from natural images. Therefore, in fig. 5, we present the results generated under the influence of multiple conditions, striving to minimize conflicts between these conditions. Utilizing textual guidance, CRS-Diff excels in generative capacity, controllability, and realism, successfully completing the synthesis of the target image. Despite the weak conditional information of our image (on the left side of the third row), a clear result is still achievable. However, without textual guidance, the quality of the synthesized image significantly degrades (the last two rows), evident in blurred lines and indistinct borders.
Table 2: Quantitative results comparisons among various CLIP models. The best results are in bold
Method IS $\uparrow$ FID $\downarrow$ CLIP Score ($\times$100) $\uparrow$ Zero-Shot classification OA $\uparrow$
CRS-Diff+CLIP(ViT-L) 14.96 108.01 19.89 53.40
CRS-Diff+CLIP(ViT-L*) 16.34 71.88 19.98 65.14
CRS-Diff+CLIP(ViT-bigG) 17.68 63.68 19.99 58.17
CRS-Diff+CLIP(ViT-bigG*) 18.39 50.72 20.33 67.38
Note: * Represents models fine-tuned on remote sensing data.
4.5 Ablation Analysis
For ablation experiments, our study discusses the Single image generation task mainly on the RSICD dataset, and we analyse the impact of CRS-Diff generation across various CLIP model bootstrapping approaches in table 2. The results show that both choosing a model with a larger number of parameters and fine-tuning on remote sensing images contribute significantly to the performance.
Table 3: Experimental results comparison of road detection in the setting of different scaled datasets. Both the real and synthetic datasets consist of 5000 images. The best results are in bold
[…]
Real 48.69 57.40 82.73 67.77
Real + Synthetic 48.98 58.89 81.75 67.92
4.6 Augmentation by our generated images
For the condition generation phase, we posit that the generated image should have a sufficiently high correlation with the control labels as conditions. We aim for the generated image to encapsulate as much information from the conditioned image as possible and to offer data support for downstream tasks in related fields. Consequently, we incorporate experiments on the generated images to ascertain the pertinent performance of CRS-Diff. We consistently integrate synthetic dummy data into the training set of SGCN [51] for the road extraction task and assess it on the official test set. Experimental results in table 3 verify that the synthetic data can improve the performance of SGCN to some extent, which means that the generated RS images conditioned by road are beneficial to downstream road detection task.
5 Conclusion
In this paper, we propose a new controllable RS image generation model (CRS-Diff), rooted in the foundational diffusion model architecture. Developed from the diffusion model framework, CRS-Diff enables high-quality remote sensing image generation. By integrating an optimized multi-conditional control mechanism, CRS-Diff can effectively synthesize multidimensional information, including text and images, guiding precise image generation and yielding highly accurate and controllable remote sensing images. This significantly broadens the control spectrum of the generation model, thus improving its adaptability to more complex application scenarios. Additionally, the generated images can help improve the performance of road detection task.
Supplementary Materials
Appendix 0.A Additional Training Details
We will detail the experimental process in separate sections, mainly including the two training phases, the datasets, and the corresponding processing utilized during training.
Single image generation. we utilized the corresponding weights from the Diffusers library and SD1.5, employing solely the RSICD dataset. Wherein during training, we randomly selected between 1 to 5 text descriptions for training data pairs. During the inference process, we generated 10 synthetic images for each text pair (excluding the untagged part of the original dataset), subsequently training a ResNet-18 model for evaluating the experimental results of IS, FID, according to the experimental criteria of Txt2Img. In zero-shot classification, we performed zero-sample classification on 2054 real images in the original test set and obtained the corresponding OA scores.
Compositional image generation. we initially reconstructed the relevant datasets (fMoW dataset and Million-AID dataset) to prepare for the subsequent experimental setup. Notably, in the caption field of the fMoW dataset, we added two easily comprehensible fields (Month and Cloud cover) to the text description and discarded them at a 20% probability. For the metadata, we encoded them into vectors of uniform length, utilizing the same embedding layer structure as the time step, and subsequently aggregated them to offer guidance. It is crucial to note that the impact of the control exerted by individual metadata varies significantly, with some fields, such as month, exhibiting a pronounced effect, while others display a weak influence, aligning with objective realities.
Appendix 0.B Additional Samples
0.B.1 Single Image Generation
Figure 6: Illustration of diversity results for single text generation of our method. Eight images are randomly generated corresponding to each given text description. No active effort is made to select visually superior synthetic images; the aforementioned images are all randomly generated from the model.
0.B.2 Compositional Image Generation
Figure 7: Illustration of diversity results under the control of a single condition of our method. Eight images are randomly generated corresponding to each given text description.
Figure 8: Illustration of diversity results for multiple conditions of our method. Eight images are randomly generated corresponding to each given text description.